Earth and Artistry, Inc. provides commercial landscaping services. Sasha Cairns, the firm’s owner, wants to develop cost estimates that she can use to prepare bids on jobs. After analyzing the firm’s costs, Cairns has developed the following preliminary cost estimates for each 1,000 square feet of landscaping.

Direct material $ 440
Direct labor (5 direct-labor hours at $10 per hour) 50
Overhead (at $16 per direct-labor hour) 80
Total cost per 1,000 square feet $ 570

  

Cairns is quite certain about the estimates for direct material and direct labor. However, she is not as comfortable with the overhead estimate. The estimate for overhead is based on the overhead costs that were incurred during the past 12 months as presented in the following schedule. The estimate of $16 per direct-labor hour was determined by dividing the total overhead costs for the 12-month period ($632,000) by the total direct-labor hours (39,500).

Regular Direct- Overtime Direct- Total Direct-
Total Overhead Labor Hours Labor Hours* Labor Hours
January $ 56,000 3,460 160 3,620
February 50,000 3,380 30 3,410
March 47,000 3,380 60 3,440
April 53,000 2,460 220 2,680
May 55,000 3,030 400 3,430
June 55,000 2,860 770 3,630
July 49,000 2,450 670 3,120
August 47,000 2,710 330 3,040
September 57,000 3,170 70 3,240
October 54,000 3,450 40 3,490
November 60,000 3,120 40 3,160
December 49,000 2,980 260 3,240
Total $ 632,000 36,450 3,050 39,500

*The overtime premium is 50 percent of the direct-labor wage rate.

Cairns believes that overhead is affected by total monthly direct-labor hours. Cairns decided to perform a least-squares regression of overhead (OH) on total direct-labor hours (DLH). The following regression formula was obtained.

OH = $43,822 + $2.69 DLH

Required:

  1. Using the overhead formula that was derived from the least-squares regression, determine a total variable-cost estimate for each 1,000 square feet of landscaping.

  2. Cairns has been asked to submit a bid on a landscaping project for the city government consisting of 56,000 square feet. Cairns estimates that 60 percent of the direct-labor hours required for the project will be on overtime. Calculate the incremental costs that should be included in any bid that Cairns would submit on this project. Use the overhead formula derived from the least-squares regression.

  1. After attending a seminar on activity-based costing, Cairns decided to further analyze the company’s activities and costs. She discovered that a more accurate portrayal of the firm’s cost behavior could be achieved by dividing overhead into three separate pools, each with its own cost driver. Separate regression equations were estimated for each of the cost pools, with the following results.

OH1 = $10,700 + $4.00 DLH,

where DLH denotes direct-labor hours

OH2= $9,100 + $13.50 SFS,

where SFS denotes the number of square feet of turf seeded (in thousands)

OH3 = $8,600 + $6.30 PL,

where PL denotes the number of individual plantings (e.g., trees and shrubs)

Assume that 5 direct-labor hours will be needed to landscape each 1,000 square feet, regardless of the specific planting material used.

     a. Suppose the landscaping project for the city will involve seeding all 56,000 square feet of turf and planting 90 trees and shrubs. Calculate the incremental variable overhead cost that Cairns should include in the bid.

     b. Recompute the incremental variable overhead cost for the city’s landscaping project assuming half of the 56,000-square-foot landscaping area will be seeded and there will be 250 individual plantings. The plantings will cover the entire 56,000-square-foot area.
Answer #1
1 Regression formula:
OH = $43822 + $2.69 DLH
Variable overhead cost=$ 2.69 per DLH
Total variable cost:
$
Direct material 440
Direct labor (5*10) 50
Variable overhead (5*2.69) 13.45
Total 503.45
2 Landscape in thousands of square feet=56000/1000=56
Incremental cost:
$
Direct material (56*440) 24640
Direct labor (56*50) 2800
Variable overhead (56*2.69) 150.64
Overtime (Note:1) 2100
Total 29690.64
Note:1
Overtime in hours=Total DLH*60%=56*5*50%=140 DLH
Overtime rate per hour=DL wage rate+(DL wage rate*50%)=10+(10*50%)=10+5=$ 15
Overtime cost=Overtime in hours*Overtime rate per hour=140*15=$ 2100
3
a. Incremental variable overhead cost:
$
OH1 (56*5*4) 1120
(Landscape in 1000 sq ft*DLH per 1000 sq ft*4)
OH2 (56*13.50) 756
(Landscape in 1000 sq ft*13.50)
OH3 (90*6.30) 567
(Number of individual plantings*6.30)
Total 2443
b. Number of square feet of turf seeded (in thousands)=(56000*50%)/1000=28
Incremental variable overhead cost:
$
OH1 (28*5*4) 560
(Landscape in 1000 sq ft*DLH per 1000 sq ft*4)
OH2 (28*13.50) 378
(Landscape in 1000 sq ft*13.50)
OH3 (250*6.30) 1575
(Number of individual plantings*6.30)
Total 2513